OpenStudy (anonymous):

Use the following table to answer the question. X A B C D P(X) 0.30 0.20 ? 0.28 Find P(B and D) assuming that B and D are independent.

OpenStudy (amorfide):

well first of you need to work out the ? p(x) is the probability they add up to 1 so add up 0.3 + 0.2 + 0.28=0.78 1-0.78=0.22 C's probability is 0.22 i can not help any further sorry :/ i need to know this too lol

OpenStudy (anonymous):

thank you for trying

OpenStudy (ash2326):

I'll take it from here :)

OpenStudy (ash2326):

We need to find , we found P(C) as 0.22 P(B and D) ?? If B and D are independent, which means occurrence of B doesn't have any effect on the occurrence of C and viceversa so \[P(B\ and\ D)= P(B)\times P(D)\] I think you can do now

OpenStudy (anonymous):

would it be .056??

OpenStudy (ash2326):

Yeah:) you're right

OpenStudy (anonymous):

yeahhhh :)

OpenStudy (amorfide):

how? lol

OpenStudy (ash2326):

@amorfide If two events are independent then the probability of their simultaneous occurrence is equal to the product of their probabilities

OpenStudy (amorfide):

nevermind i multiplied wrong numbers :$